### v1

@@ 5.4.2 Hyperparameters
Table 6: Hyperparameter values for the heterogeneous firms and households model.
5.5 Accuracy
In this section, we demonstrate the accuracy of our solution. Since a closed-form or a high-quality numerical reference solution is not available, we investigate the errors in the equilibrium conditions, implied by the price and policy functions encoded in the learned parameters of our neural networks. We report the accuracy measures in economically interpretable units. In addition, we provide a comprehensive set of statistics on the distribution of errors across the state space.
5.5.1 Firm policies
The optimal firm policies are characterized by the set of firms’ KKT conditions, given in equations (50) to (53). Rearranging equation (50) we obtain:
where the numerator in the first term denotes the expected marginal benefit of additional capital in the next period and the denominator gives the marginal cost in the current period. The errors in equation (112) are therefore interpretable as relative marginal cost errors. [^49] An error of 0.01, for example, would indicate that the marginal benefit is 1% higher than the marginal cost. The left panel of figure 5 shows statistics of the distribution of errors in equation (112) for different idiosyncratic productivity shocks, idiosyncratic levels of capital, and 1024 aggregate states drawn from the ergodic distribution of the economy.
The mean marginal cost error remain below 0.07%, the 90th percentile below 0.16% and the 99.9th percentile below 0.4% for all levels of firm capital and the three productivity levels. Even though the errors are very low everywhere, they are slightly larger in the areas of the state space where the firm policies feature a kink due to the constraint on dividends and due to asymmetric adjustment costs. [^50] For a fixed idiosyncratic level of capital, and a fixed idiosyncratic productivity level, the mean and the percentiles are computed across 1024 aggregate states drawn from the ergodic distribution of the economy.
where $\psi^{\text{FB}}(x,y)$ denotes the Fischer-Burmeister function. Errors in equation (113) summarize violations of the all remaining KKT conditions associated with the firm problem. [^51] Because the Fischer-Burmeister function satisfies $\psi^{\text{FB}}(x,y)=0\Leftrightarrow x\geq 0,y\geq 0,xy=0$. Except for very low levels of capital, the error is virtually zero. Throughout, the mean error remains below 0.2% and even the 99.9th percentile of errors remains below 0.8%. We conclude that the firm policies are computed with high accuracy across the aggregate and idiosyncratic state space.
5.5.2 Household policies
5.5.3 Stock price
The left panel of figure 7 shows the distribution of discrepancies between the equity price predicted by the neural network relative and the market clearing equity price obtained using the EGM-Newton-Raphson algorithm. The mean error of the price prediction is 0.17%, the 90th percentile 0.35% and the 99.9th percentile is 0.80%. The right panel of figure 7 shows the price predictions by the neural network together with the computed market clearing prices in a scatter plot against aggregate dividends.
While the accuracy of the price function is slightly lower than the accuracy we achieved for the policy functions of firms and households, it remains high. The error in the asset demand, as implied by the predicted price and the predicted policy functions is lower and below 0.1% over the whole simulated ergodic set.
5.6 Inspecting the learned equilibrium policies
5.6.1 Firms
We can see that the distribution of firms across capital varies substantially. Due to the decreasing returns to scale production function, together with capital adjustment costs, this variation directly affects aggregate wages and dividends (see, e.g., equation (59)), such that the households problem depends crucially on the firm distribution.
The middle panel in figure 9 shows the dividend policies of firms for different capital levels (horizontal axis) and different idiosyncratic productivity levels (represented by different colors). The shaded lines show the policies for several randomly selected aggregate states, and the solid lines show the mean taken over aggregate states. Each of the policy functions shows two kinks: at the first kink, the non-negativity constraint on dividends stops to bind. At the second kink, firms switch from adjusting capital upward to adjusting it downward. The exact location of the kinks depends on the productivity level and the firms capital as well as on the aggregate state of the economy.
Similarly, the right panel shows the policy functions for the KKT multiplier on the dividend constraint. Perhaps surprisingly, the multiplier for very low values of capital is larger for the low productivity values than for the high productivity values. This is the case because, while the firms are constraint for all three productivity levels, the high productivity firms are able to choose a higher capital level for the next period.
5.6.2 Households
Figure 9: Left and middle panel: consumption function of households for their idiosyncratic productivity level (in different colors) and for different idiosyncratic asset holdings (horizontal axis). The shaded lines show the consumption policies for several different aggregate states, and the solid lines show the mean across aggregate states. Right panel: marginal propensity to consume for different productivity and wealth levels, as well as for several aggregate states.
As figure 9 shows, consumption varies substantially between different aggregate states. We can also see that, for each of the aggregate states, consumption is increasing and concave in the wealth of households. Our shape preserving neural network architecture ensures that this economically important feature of the consumption functions is guaranteed. We achieve this using the three step procedure described above. The households’ marginal propensities to consume are constructed from the predictions of the neural network, such that the marginal propensity to consume, for a given aggregate state and idiosyncratic productivity level, is always decreasing and convex in cash-at-hand, as well as bounded between 1 and 0 (see the right panel in figure 9). This ensures on the one hand the monotonicity and concavity of the consumption function, and on the other hand facilitates the precise prediction of the marginal propensity to consume (and hence consumption) of borrowing constraint households.
6 Conclusion
We introduce a new algorithm for computing global solutions of dynamic stochastic general equilibrium models. Exploiting the ergodic property of a large class of dynamic economies, we rely on the truncated history of aggregate shocks as an approximate sufficient statistic for the aggregate state. Based on this approximation, we use deep neural networks to parameterize the mapping from truncated aggregate shock histories to equilibrium objects of interest using deep neural networks. Finally, we train our neural networks to satisfy all the equilibrium conditions along simulated paths of the economy.
We illustrate the accuracy and wide applicability of our algorithm by solving three example economies. As a first example and proof of concept, we solve a simple stochastic growth model. This allows us to compare our solution against a high-accuracy solution that can be obtained using conventional grid-based methods. Second, we solve an overlapping generations model with 72 age groups, portfolio choice, and multiple sources of aggregate risk. The overlapping generations model shows that the sequence space approach is applicable, also when there is a clear dependence of equilibrium objects on long histories of shocks: the shocks experienced during an agent’s lifetime affect their subsequent equilibrium outcomes. The model includes stochastic shocks to productivity and depreciation, which are drawn from normal distributions, together with a Markov regime switching model with two discrete states. We thereby illustrate that our method is capable of solving models with several aggregate shocks of different types. As a third example, we consider an economy where a continuum of heterogeneous households trade in a long-lived financial asset, which constitutes a claim to the dividends paid by a continuum of heterogeneous firms, who operate a decreasing returns to scale production technology and face asymmetric adjustment costs on capital. Additionally, the economy is subject to stochastic fluctuations in aggregate productivity and in the level of aggregate and idiosyncratic uncertainty. The state hence includes two endogenously moving distributions: the distribution over wealth and productivity on the household side and the distribution of firms over productivity and capital. We show that we solve all three models accurately, with mean relative errors in equilibrium conditions below 0.1%.
To solve models with aggregate and idiosyncratic risk, we introduce shape preserving operator learning: we train deep neural networks to predict idiosyncratic policy functions as a function of the truncated history of aggregate shocks, such that we can guarantee the predicted policy functions to be monotone or concave in key idiosyncratic state variables, such as household wealth. These guaranties allow us to use the method of endogenous gridpoints and a simple, yet-reliable, Newton-Raphson algorithm to obtain high-quality training targets for supervised learning of household policies and the market clearing stock price, further increasing the robustness of our deep learning algorithm.

### v2

@@ Appendix A Additional Tables
Table A6: Hyperparameter values for the heterogeneous firms and households model.
Parameter Meaning Value
$H$ Number of cohorts $10$
$\gamma$ Relative risk aversion 22
$\beta$ Patience (per 6-year period) $0.95^{6}\approx 0.735$
$\alpha$ Capital share in production $1/3$
$\underline{k}$ Borrowing constraint 0
$N_{k}$ Asset grid points $300$
$k\in[\underline{k},\bar{k}]$ Asset grid bounds $[0,5]$
$\mathcal{E}^{\text{grid}}$ Idiosyncratic productivity states $\{0.5,1.0,1.5\}$
$\Pi^{e}$ Transition matrix for $e$ See text
$\delta_{n}$ Depreciation (normal) $1-(1-0.05)^{6}\approx 0.265$
$\delta_{d}$ Depreciation (disaster) $1.5\delta_{n}\approx 0.397$
$\Pi^{z}$ Transition matrix for $z$ See text
$\epsilon$ Utility cost of work at $h=6$ $1.68$
$\sigma$ Taste-shock scale (logit smoothing) $0.08$
$y^{\text{home}}$ Income when retired $0.1$
$\psi^{b}$ Bequest motive strength $0.1$
$\nu$ Bequest curvature 22
$\bar{b}$ Bequest shifter $0.1$
Table A7: Parameter values for the discrete-continuous choice OLG model in section 6.
Appendix B Additional Figures
Figure A1: Efficiency units of labor over the life-cycle for the OLG model in section 4.
Figure A2: Continuation values by age and asset holdings for different idiosyncratic productivity and several aggregate states.
Figure A3: The distribution of (absolute) prediction errors for the continuation value by age and asset holdings.
Figure A4: Consumption functions of retired cohorts. The shaded lines show the policies for different aggregate states, and the solid lines show the average taken across the sample of aggregate states.